# Dimension-independent convergence rates of randomized nets using median-of-means

### v3

@@ 1 Introduction
The quasi-Monte Carlo (QMC) method is a powerful alternative to conventional Monte Carlo (MC) techniques for high-dimensional integration. While both methods approximate integrals by averaging $n$ function evaluations, QMC replaces MC’s pseudorandom sampling with low-discrepancy point sets engineered to uniformly explore the sampling space. This structured approach allows QMC to avoid the exponential cost growth with dimension of classical quadrature rules while achieving faster error convergence than MC. These advantages have made QMC particularly valuable in fields like computational finance [joy1996quasi, l2009quasi] and physics [kuo:nuye:2016], where high-dimensional integrals are pervasive.
@@
Building on these strengths, we focus on randomized quasi-Monte Carlo (RQMC) methods based on linearly scrambled base-2 digital nets. Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [dick:pill:2010]. The random linear scrambling [mato:1998:2], a computationally efficient alternative to Owen’s scrambling [rtms], preserves the structure of digital nets while introducing controlled randomness. This randomization enhances the root mean squared error convergence rate to nearly $O(n^{-3/2})$ under smoothness conditions and enables statistical inference. However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [superpolyone].
@@
To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [chen2025randomintegrationalgorithmhighdimensional, goda2024simpleuniversalalgorithmhighdimensional, goda:lecu:2022, Goda2024, ye2025medianqmcmethodunbounded]. The median of linearly scrambled digital net estimators was first analyzed in [superpolyone], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$. Subsequent work in [superpolymulti] extended this result to higher dimensions. For finitely differentiable integrands, [pan2024automatic] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
@@
A critical gap in these prior studies is the role of dimensionality. We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [kuo:sloa:wasi:wozn:2010]. Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [nova:wozn:2008], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3). This notion of relative variation aligns with the effective dimension framework introduced in [cafmowen], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions. The connection to weighted function spaces [sloa:wozn:1998] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
@@
Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [Goda2024]. A key distinction lies in the error criterion: their analysis employs the worst-case integration error over the unit ball in a weighted Sobolev space, whereas we demonstrate that median estimators achieve near-optimal convergence rates on a fixed integrand with high probability. Our probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [kunsch2019solvable, kunsch2019optimal].
@@
Given $\eta>0$, $k\in\mathbb{N}$ and $f\in C^{(k,\dots,k)}([0,1]^{s})$, it holds with probability $1-2^{-r}$ that the median estimator $\hat{\mu}^{(r)}_{\infty}$ from Subsection 2.3 satisfies
$|\hat{\mu}^{(r)}_{\infty}-\mu|\leqslant C_{\eta,k,f}n^{-k-1/2+\eta}.$
The constant $C_{\eta,k,f}$ is independent of $s$ under the random linear scrambling if the relative variation $\gamma_{u}$ satisfies the tractability conditions in Theorem 5.
These nearly $O(n^{-k-1/2})$ integrand-specific error rates complement the nearly $O(n^{-k})$ worst-case rates established in [Goda2024]. Furthermore, Theorem 5 requires strictly weaker assumptions than those assumed in [Goda2024] for the random linear scrambling, demonstrating that dimension-independent convergence is more easily achieved under our error criterion.
As a final remark, we note that while several algorithms match lower complexity bounds for integration in specific spaces—such as higher-order digital nets [goda2016explicit, goda2017optimal, goda2018optimal] and Frolov lattice rules [frolov1976upper, krieg2017universal, ullrich2016role]—our method offers distinct practical advantages. It requires no smoothness parameter specification, yet automatically matches theoretical lower bounds up to an arbitrarily small $\eta>0$ in the convergence rate, making it both efficient and immune to parameter misspecification. Combined with its straightforward implementation and the dimension-independent convergence established in this paper, the median estimator presents a highly practical and robust alternative for applied problems. Furthermore, recent work has successfully extended median RQMC to $L^{2}$-approximation [pan2025universal, pan2025l_2], demonstrating its broader utility beyond numerical integration.
@@ 2.1 Digital net construction and randomization
where $C_{j}\in\{0,1\}^{E\times m}$ and $\vec{D}_{j}\in\{0,1\}^{E}$ are random with precision $E\geqslant m$. The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli(0.5)(0.5) entries. A common way to randomize $C_{j}$ is the random linear scrambling [mato:1998:2]:
@@
where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli(0.5)(0.5)), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [dick:pill:2010, Chapter 4.4] for details).
While effective, the reliance of random linear scrambling on carefully designed generating matrices complicates theoretical analysis. We address this in Section 5. For initial insights, Section 4 adopts the complete random design [pan2024automatic], where all entries of $C_{j}$ are drawn independently from Bernoulli(0.5)(0.5). This approach retains the asymptotic convergence rate of random linear scrambling while circumventing the need for pre-optimized generator matrices.
@@
We conveniently assume $E=\infty$ and focus our theoretical analysis on $\hat{\mu}_{\infty}$. In practice, $E$ is constrained by computational precision. Lemma 1 of [superpolymulti] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$. Thus, $\hat{\mu}_{E}$ closely approximates $\hat{\mu}_{\infty}$ when $\omega_{f}(\sqrt{s}2^{-E})\ll|\hat{\mu}_{E}-\mu|$.
@@ 2.2 Walsh functions and error analysis for base-2 digital nets
where $\boldsymbol{k}=(k_{1},\dots,k_{s})\in\mathbb{N}^{s}_{0}$. These functions constitute a complete orthonormal basis for $L^{2}([0,1]^{s})$ [dick:pill:2010], enabling the Walsh series decomposition:
@@
The equality holds in the $L^{2}$-sense. Using this decomposition, [pan2024automatic, Theorem 1] establishes a key error structure for QMC estimators:
@@ 2.3 Probabilistic error guarantee of the median estimator
First note that the random signs $S(\boldsymbol{k})$ in equation (4) are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [superpolyone, Lemma 4] for the one-dimensional case). Since $\mathcal{A}$ is independent of $S(\boldsymbol{k})$, we have $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and
@@ 2.4 Walsh coefficients and function smoothness
Equation (5) comes from [SUZUKI20161, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [SUZUKI20161, Section 3]. ∎
For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$. In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3. This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (5), as discussed in [pan2024automatic, Remark 1].
@@ 3 ANOVA decomposition and effective dimensionality
To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [meandim]:
@@
To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [dick:2008], which generalizes the usual Sobolev norms. For a function $f:[0,1]^{d}\to\mathbb{R}$ and a subcube $J=\prod_{j=1}^{d}[a_{j},b_{j})$, we define the alternating sum over $J$:
@@
with equality attained when $f\in C^{(1,\dots,1)}([0,1]^{s})$ [dick:2008, equation (3.4)]. More generally, [LIU2025101935] shows $V^{(d)}_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$. Discontinuous $f$ can also have finite $V^{(d)}_{\lambda}(f)$ (see [pan2024automatic, Section 4.1] for an example).
@@
By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [pan2024automatic, Section 3.2],
@@ 4 Tractability under the completely random design
with the sum from 11 to $\alpha$ interpreted as 0 when $\alpha=0$.
@@
where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is nearly optimal, as it aligns with the lower bound established in [bakh:1959] for one-dimensional $f$ with $\lambda$-Hölder continuous $f^{(\alpha)}$, a subclass of functions with finite $\|f\|_{\alpha,\lambda}$.
@@
which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [HICKERNELL2003286, Lemma 3]. Putting the above bounds into equation (22), we conclude that equation (24) holds with a constant $C_{\eta}$ independent of $s$. Notably, equation (25) resembles the tractability conditions assumed by [Goda2024, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
@@ 5 Tractability under more general randomization
While the completely random design offers theoretical advantages in near-optimal convergence rates and tractability, the random linear scrambling often outperforms it in practical settings due to tighter error bounds on low-dimensional ANOVA components. For example, [joe:kuo:2008] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets. In this section, we extend the analysis of Section 4 to a broader class of randomization schemes, with the random linear scrambling as a key example.
@@
From [Goda2024, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [Goda2024, Section 2]. We frame this assumption to align with broader applications in higher-order digital nets [dick:2011], where the parameter $d$ (the order of digital nets) may exceed $1$. This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
with the sum from 11 to $\alpha-d$ interpreted as 0 when $\alpha=d$.
@@
Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [c812a4a5-75ac-3614-936d-5782e44941d0] that the base-2 Niederreiter sequence [NIEDERREITER198851] satisfies
@@
while the Sobol’ sequence [sobol67] satisfies
@@
While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [dick:2011], and unbounded increases in $d$ ultimately violate the tractability condition above. Thus, prudent selection of $d$ is critical for practical implementations.
@@
Since $\phi>1$, this condition is strictly weaker than the requirements in [Goda2024, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates under the random linear scrambling.

### v4

@@ 1 Introduction
The quasi-Monte Carlo (QMC) method is a powerful alternative to conventional Monte Carlo (MC) techniques for high-dimensional integration. While both methods approximate integrals by averaging $n$ function evaluations, QMC replaces MC’s pseudorandom sampling with low-discrepancy point sets engineered to uniformly explore the sampling space. This structured approach allows QMC to avoid the exponential cost growth with dimension of classical quadrature rules while achieving faster error convergence than MC. These advantages have made QMC particularly valuable in fields like computational finance [19, 25] and physics [23], where high-dimensional integrals are pervasive.
@@
Building on these strengths, we focus on randomized quasi-Monte Carlo (RQMC) methods based on linearly scrambled base-2 digital nets. Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [4]. The random linear scrambling [28], a computationally efficient alternative to Owen’s scrambling [31], preserves the structure of digital nets while introducing controlled randomness. This randomization enhances the root mean squared error convergence rate to nearly $O(n^{-3/2})$ under smoothness conditions and enables statistical inference. However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [34].
@@
To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [3, 10, 11, 12, 44]. The median of linearly scrambled digital net estimators was first analyzed in [34], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$. Subsequent work in [35] extended this result to higher dimensions. For finitely differentiable integrands, [36] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
@@
A critical gap in these prior studies is the role of dimensionality. We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [24]. Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [30], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3). This notion of relative variation aligns with the effective dimension framework introduced in [2], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions. The connection to weighted function spaces [38] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
@@
Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [12]. A key distinction lies in the error criterion: their analysis employs the worst-case integration error over the unit ball in a weighted Sobolev space, whereas we demonstrate that median estimators achieve near-optimal convergence rates on a fixed integrand with high probability. Our probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [21, 22].
@@
Given $\eta>0$, $k\in\mathbb{N}$, $r\in\mathbb{N}$ and $f\in C^{(k,\dots,k)}([0,1]^{s})$, it holds with probability $1-2^{-r}$ that the median estimator $\hat{\mu}^{(r)}_{\infty}$ from Subsection 2.3 satisfies
$\left|\hat{\mu}^{(r)}_{\infty}-\int_{[0,1]^{s}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}\right|\leqslant C_{\eta,k,f}n^{-k-1/2+\eta}.$
Here, $C^{(k,\dots,k)}([0,1]^{s})$ denotes functions with continuous dominating mixed partial derivatives of order $k$ (see Subsection 2.4). The constant $C_{\eta,k,f}$ is independent of $s$ under the random linear scrambling if the relative variation $\gamma_{u}$ satisfies the tractability conditions in Theorem 5.
These nearly $O(n^{-k-1/2})$ integrand-specific error rates complement the nearly $O(n^{-k})$ worst-case rates established in [12]. Furthermore, Theorem 5 requires strictly weaker assumptions than those assumed in [12] for the random linear scrambling, demonstrating that dimension-independent convergence is more easily achieved under our error criterion.
As a final remark, we note that while several algorithms match lower complexity bounds for integration in specific spaces—such as higher-order digital nets [13, 14, 15], Frolov lattice rules [8, 20, 42] and Smolyak algorithms [9, 37, 39]—our method offers distinct practical advantages. It requires no smoothness parameter specification, yet automatically matches theoretical lower bounds up to an arbitrarily small $\eta>0$ in the convergence rate, making it both efficient and immune to parameter misspecification. Combined with its straightforward implementation and the dimension-independent convergence established in this paper, the median estimator presents a highly practical and robust alternative for applied problems. Furthermore, recent work has successfully extended median RQMC to $L^{2}$-approximation [32, 33], demonstrating its broader utility beyond numerical integration.
@@ 2.1 Digital net construction and randomization
where $C_{j}\in\{0,1\}^{E\times m}$ and $\vec{D}_{j}\in\{0,1\}^{E}$ are random with precision $E\geqslant m$. The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli(0.5)(0.5) entries. A common way to randomize $C_{j}$ is the random linear scrambling [28]:
@@
where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli(0.5)(0.5)), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [4, Chapter 4.4] for details).
While effective, the reliance of random linear scrambling on carefully designed generating matrices complicates theoretical analysis. We address this in Section 5. For initial insights, Section 4 adopts the complete random design [36], where all entries of $C_{j}$ are drawn independently from Bernoulli(0.5)(0.5). This approach retains the asymptotic convergence rate of random linear scrambling while circumventing the need for pre-optimized generator matrices.
@@
We conveniently assume $E=\infty$ and focus our theoretical analysis on $\hat{\mu}_{\infty}$. In practice, $E$ is constrained by computational precision. Lemma 1 of [35] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$. Thus, $\hat{\mu}_{E}$ closely approximates $\hat{\mu}_{\infty}$ when $\omega_{f}(\sqrt{s}2^{-E})\ll|\hat{\mu}_{E}-\mu|$.
@@ 2.2 Walsh functions and error analysis for base-2 digital nets
where $\boldsymbol{k}=(k_{1},\dots,k_{s})\in\mathbb{N}^{s}_{0}$. These functions constitute a complete orthonormal basis for $L^{2}([0,1]^{s})$ [4], enabling the Walsh series decomposition:
@@
The equality holds in the $L^{2}$-sense. Using this decomposition, [36, Theorem 1] establishes a key error structure for QMC estimators:
@@ 2.3 Probabilistic error guarantee of the median estimator
First note that the random signs $S(\boldsymbol{k})$ in equation (4) are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [34, Lemma 4] for the one-dimensional case). Since $\mathcal{A}$ is independent of $S(\boldsymbol{k})$, we have $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and
@@ 2.4 Walsh coefficients and function smoothness
Equation (5) comes from [41, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [41, Section 3]. ∎
For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$. In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3. This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (5), as discussed in [36, Remark 1].
@@ 3 ANOVA decomposition and effective dimensionality
To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [26]:
@@
To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [5], which generalizes the usual Sobolev norms. For a function $f:[0,1]^{d}\to\mathbb{R}$ and a subcube $J=\prod_{j=1}^{d}[a_{j},b_{j})$, we define the alternating sum over $J$:
@@
with equality attained when $f\in C^{(1,\dots,1)}([0,1]^{s})$ [5, equation (3.4)]. More generally, [27] shows $V^{(d)}_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$. Discontinuous $f$ can also have finite $V^{(d)}_{\lambda}(f)$ (see [36, Section 4.1] for an example).
@@
By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [36, Section 3.2],
@@ 4 Tractability under the completely random design
with the sum from $1$ to $\alpha$ interpreted as 0 when $\alpha=0$.
@@
where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is nearly optimal, as it aligns with the lower bound established in [1] for one-dimensional $f$ with $\lambda$-Hölder continuous $f^{(\alpha)}$, a subclass of functions with finite $\|f\|_{\alpha,\lambda}$.
@@
which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [17, Lemma 3]. Putting the above bounds into equation (22), we conclude that equation (24) holds with a constant $C_{\eta}$ independent of $s$. Notably, equation (25) resembles the tractability conditions assumed by [12, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
@@ 5 Tractability under more general randomization
While the completely random design offers theoretical advantages in near-optimal convergence rates and tractability, the random linear scrambling often outperforms it in practical settings due to tighter error bounds on low-dimensional ANOVA components. For example, [18] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets. In this section, we extend the analysis of Section 4 to a broader class of randomization schemes, with the random linear scrambling as a key example.
@@
From [12, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [12, Section 2]. We frame this assumption to align with broader applications in higher-order digital nets [6], where the parameter $d$ (the order of digital nets) may exceed $1$. This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
with the sum from $1$ to $\alpha-d$ interpreted as 0 when $\alpha=d$.
@@
Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [43] that the base-2 Niederreiter sequence [29] satisfies
@@
while the Sobol’ sequence [40] satisfies
@@
While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [6], and unbounded increases in $d$ ultimately violate the tractability condition above. Thus, prudent selection of $d$ is critical for practical implementations.
@@
Since $\phi>1$, this condition is strictly weaker than the requirements in [12, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates under the random linear scrambling.